Water is contained in a rigid vessel of 5 m3 at a quality of 0.8 and a pressure of 2 MPa. If the a pressure is reduced to 400 kPa by cooling the vessel, find the final mass of vapor mg and mass of liquid mf.
1
Expert's answer
2018-05-29T10:58:07-0400
v= vf +x (vg – vf)
v= 0.001084 + 0.8 (0.46242 – 0.001084)
v= 0.3732104 m3 / kg
v =V/m → m = V/v
m= 5 m3 / 0.3732104 m3 / kg = 13.4 kg
mg = mx = 13.4 kg x 0.8 = 10.72 kg
mf = m - mg = 13.4 kg - 10.72 kg = 2.68 kg
Answer: 10.72 kg and 2.68 kg